# FinTagging: An LLM-ready Benchmark for Extracting and Structuring Financial Information

### v1

@@ Appendix E Evaluation Models Details
  • Closed-source LLMs: We include GPT-4o [11], accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks. Although model size details are undisclosed, GPT-4o serves as an upper-bound reference in our benchmark.
  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) [13], DeepSeek-R1-Distill-Qwen (32B) [6], and multiple variants of Qwen2.5-Instruct  [34](ranging from 0.5B to 14B). We also include LLaMA-3.2 and 3.1 variants  [9] (3B and 8B), as well as Google’s Gemma-2-27B  [26], to ensure architectural diversity and scalability comparison. These models are primarily instruction-tuned and optimized for general-purpose NLP tasks.
  • Financial-specific LLMs: We evaluate Fino1-8B  [19], a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures. This category allows us to assess the benefits of domain adaptation in complex tagging and reasoning tasks.
  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  [7], FinBERT  [1], and SECBERT  [14]. These models have been widely used in prior financial NLP tasks and allow for a comparative analysis between generative LLMs and traditional pretrained models in terms of domain understanding and structured output capability.
@@
Table 10: Model Categories and Corresponding Repositories
@@
Qwen2.5-Instruct 0.5B, 1.5B, 14B Qwen/Qwen2.5-*B-Instruct
Llama-3.2-Instruct 3B meta-llama/Llama-3.2-3B-Instruct
Llama-3.1-Instruct 8B meta-llama/Llama-3.1-8B-Instruct
@@
Appendix F The details for the fine-tuning PTMs
F.1 Training data collection and processing
Similar to the collection process for the FinTagging benchmark data, we gathered an additional 10 annual 10-K financial reports filed with the SEC for the period from February 13, 2024, to February 13, 2025, as summarized in Table 11. These reports contain a total of 33,848 standard taxonomy-tagged facts. Using BeautifulSoup to parse these documents, we identified 22,847 narrative sentences (approximately 5.5 million characters) and 1,236 financial tables. The companies included in this dataset follow the XBRL standard, ensuring comprehensive coverage for training PTMs.
Table 11: Financial report statistics summary for raw training data
@@ F.1 Training data collection and processing
Table 12: The statistical information for training data. Using the “cl100k_base” tokenizer to calculate tokens ($\pm$ standard deviation).
@@
However, to align with the extreme classification format used in previous XBRL tagging benchmarks, we directly adopt the US-GAAP tags as entity labels, annotating each token in sentences and tables using the BIO scheme. Specifically, B denotes the beginning of an entity phrase, I marks the continuation (inside) of an entity phrase, and O indicates tokens outside of any entity. As shown in Figure 4, "4.9" and "4.5" are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”). To comprehensively cover all US-GAAP tags, we combine the entire set of 17,388 tags from the US-GAAP 2024 taxonomy with the BIO labeling scheme to construct an extreme classification label space, resulting in 34,777 unique entity labels ($2\times 17388+1$).
Figure 4: An example for training set annotation with BIO scheme.
@@
F.2 Training settings
We fine-tune three pretrained models, BERT-large [7], FinBERT [1], and SECBERT [14], on our training set using the HuggingFace Transformers library. All models are trained with a batch size of 4, a learning rate of 3e-5, and for 20 epochs. Optimization is performed using AdamW without gradient accumulation or early stopping. Token classification heads are randomly initialized and trained jointly with the base encoder. Input sequences are tokenized with a maximum length of 512, and labels are aligned at the sub-token level following the BIO tagging scheme. Loss is computed only on the first sub-token of each word to avoid misalignment bias.
@@ F.2 Training settings
Appendix G Retrieval Results in FinCL subtask
@@
Table 13: The Acc@k performance of retrieval for different context construction strategies on the FinCL task. Here, FWC denotes the fixed window-based strategy, while SAC denotes the structure-aware approach that respects sentence or row boundaries.
@@
SAC Sentence 0.0689 0.2643 0.3521 0.4711 0.5611 0.6096 0.6522
Table 0.0147 0.0867 0.1240 0.1855 0.2353 0.2632 0.2866
Overall 0.0191 0.1011 0.1424 0.2086 0.2617 0.2913 0.3163
@@
From Table 13, we found the SAC strategy consistently outperforms FWC, particularly in the table context, highlighting the importance of aligning the context window with the underlying structural unit. For sentence-based entities, both strategies yield comparable results, with SAC achieving slightly higher Acc@1 (0.0689 vs. 0.0677) and showing marginal improvements across all cutoff values. This suggests that even for relatively unstructured text, preserving sentence boundaries may provide minor benefits over a fixed-length window.
@@
In contrast, for table-based entities, the performance gap is substantial. SAC achieves significantly higher retrieval accuracy (e.g., Acc@100 of 0.2353 vs. 0.0050), indicating that row-level context is far more informative than arbitrary character windows when dealing with tabular structures. FWC performs poorly in this setting, likely due to the fragmented and semantically sparse nature of partial table text.
@@
When aggregating results across both structures, SAC outperforms FWC by a wide margin at all retrieval depths (e.g., Acc@200 of 0.3163 vs. 0.0608). These results underscore the importance of structure-aware context construction, especially in scenarios where inputs span multiple formats such as sentences and tables.
Appendix H Prompt Templates
H.1 Template for FinNI subtask
@@
Prompt Template You are a financial information extraction expert specializing in identifying financial numerical entities in XBRL reports. Your task is to extract all such numerical entities from the provided text or serialized <table></table> data and classify them into one of five categories: - "integerItemType": Counts of discrete items, such as the number of employees or total transactions. - "monetaryItemType": Financial amounts expressed in currency, such as revenue, profit, or total assets. - "perShareItemType": Per-share values, such as earnings per share (EPS) or book value per share. - "sharesItemType": Counts of shares, such as outstanding shares or ownership stakes. - "percentItemType": Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol ("%"). Important Instructions: (1) Financial numerical entities are not limited to Arabic numerals (e.g., 10,000). They may also appear in word form (e.g., "ten million"), which must be correctly identified and converted into standard numerical format. (2) Not all numbers in the text should be extracted. Only those that belong to one of the five financial entity categories above should be included. Irrelevant numbers (such as phone numbers, dates, or general IDs) must be ignored. (3) If a number is followed by a magnitude term (e.g., Hundred, Thousand, Million, Billion), do not expand it into the full numerical value. * "Two hundred" -> Extract only "two", not "200". * "10.6 million" -> Extract only "10.6", not "10,600,000". (4) Standardize numerical formatting by removing currency symbols (e.g., "USD"), percentage signs ("%"), and commas (",") while preserving the numeric value. These elements must be removed to ensure consistency. (5) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>}]} Input: {text/table} Output:
H.2 Template for ablation
@@
Prompt Template You are an XBRL tagging expert specializing in annotating financial numerical facts in XBRL reports. Your task is to (1) extract all such numerical entities from the provided text or serialized <table></table> data, (2) classify them into one of five categories, and (3) assign an appropriated US-GAAP tag to each entity. Categories: - "integerItemType": Counts of discrete items, such as the number of employees or total transactions. - "monetaryItemType": Financial amounts expressed in currency, such as revenue, profit, or total assets. - "perShareItemType": Per-share values, such as earnings per share (EPS) or book value per share. - "sharesItemType": Counts of shares, such as outstanding shares or ownership stakes. - "percentItemType": Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol ("%"). US-GAAP tags: - A US-GAAP tag is a standardized semantic label used in XBRL filings to identify specific financial concepts defined by the U.S. Generally Accepted Accounting Principles (GAAP). Each tag represents a distinct accounting item and enables consistent, machine-readable financial reporting. - Examples: "us-gaap:AssetsCurrentAbstract", "us-gaap:AccruedInsuranceNoncurrent". Important Instructions: (1) Financial numerical entities are not limited to Arabic numerals (e.g., 10,000). They may also appear in word form (e.g., "ten million"), which must be correctly identified and converted into standard numerical format. (2) Not all numbers in the text should be extracted. Only those that belong to one of the five financial entity categories above should be included. Irrelevant numbers (such as phone numbers, dates, or general IDs) must be ignored. (3) If a number is followed by a magnitude term (e.g., Hundred, Thousand, Million, Billion), do not expand it into the full numerical value. * "Two hundred" -> Extract only "two", not "200". * "10.6 million" -> Extract only "10.6", not "10,600,000". (4) Standardize numerical formatting by removing currency symbols (e.g., "USD"), percentage signs ("%"), and commas (",") while preserving the numeric value. These elements must be removed to ensure consistency. (5) You should assign the most appropriate US-GAAP tag to each identified entity based on your internal understanding of the 2024 US-GAAP taxonomy. (6) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>, "Tag": <Assigned US-GAAP tag>}]} Input: {text/table} Output:
Appendix I Potential Societal Impacts of the Work
I.1 Potential Positive Societal Impacts
This work introduces FinTagging, a full-scope, structure-aware benchmark for evaluating large language models on XBRL-based financial reporting. By supporting realistic tagging scenarios grounded in US-GAAP taxonomy, this benchmark:
  • Promotes automation and transparency in financial disclosure, which can reduce manual errors and increase the reliability of corporate filings.
  • Enables more accurate, scalable tools for financial regulators, auditors, and analysts, potentially improving oversight and decision-making in global markets.
  • Encourages the development of LLMs with stronger schema-aware reasoning, a critical step toward trustworthy AI systems in high-stakes financial environments.
Through open access to annotated datasets and evaluation code, FinTagging also advances reproducible research and supports the broader community in benchmarking financial language models responsibly.
I.2 Potential Negative Societal Impacts
While FinTagging provides valuable tools for model evaluation, several risks should be noted:
  • Misuse and Misinterpretation: Benchmark performance may be misinterpreted as real-world readiness. Deployment of LLMs without sufficient regulatory safeguards could lead to incorrect tagging, misreporting, or financial misinformation.
  • Coverage Bias: The benchmark focuses on US-GAAP financial reports and may underrepresent other accounting standards or financial contexts, potentially limiting its applicability in non-U.S. jurisdictions.
  • Overreliance on Automation: Improved tagging accuracy may incentivize overautomation in financial reporting pipelines, increasing the risk of undetected errors if human oversight is reduced.
To mitigate these concerns, we recommend that future work includes broader financial contexts and integrates FinTagging into a framework with human-in-the-loop validation, fairness checks, and error auditing. Ongoing collaboration with domain experts and regulators is also essential.

### v2

@@ Appendix H Evaluation Models Details
  • Closed-source LLMs: We include GPT-4o (Hurst et al., 2024), accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks. Although model size details are undisclosed, GPT-4o serves as an upper-bound reference in our benchmark.
  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) (Liu et al., 2024), DeepSeek-R1-Distill-Qwen (32B) (DeepSeek-AI, 2025), and multiple variants of Qwen3  (Team, 2025)(ranging from 0.6B to 32B). We also include Llama-3.3, Llama-3.2, and 3.1 variants  (Grattafiori et al., 2024) (70B, 3B, and 8B), Llama-4-Scout-17B-16E-Instruct (Meta, 2025)(109B), as well as Google’s Gemma-2-27B  (Team, 2024), to ensure architectural diversity and scalability comparison. These models are primarily instruction-tuned and optimized for general-purpose NLP tasks.
  • Financial-specific LLMs: We evaluate Fino1-8B  (Qian et al., 2025), a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures. This category allows us to assess the benefits of domain adaptation in complex tagging and reasoning tasks.
  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  (Devlin et al., 2018), FinBERT  (Araci, 2019), and SECBERT  (Loukas et al., 2022). These models have been widely used in prior financial NLP tasks and allow for a comparative analysis between generative LLMs and traditional pretrained models in terms of domain understanding and structured output capability.
@@
Table 13. Model categories and corresponding repositories.
@@
Llama-4-Scout-17B-16E-Instruct 109B meta-llama/Llama-4-Scout-17B-16E-Instruct
Llama-3.3-70B-Instruct 70B meta-llama/Llama-3.3-70B-Instruct
@@
Qwen3-32B 32B Qwen/Qwen3-32B
@@
Qwen3-14B 14B Qwen/Qwen3-14B
Llama-3.1-8B-Instruct 8B meta-llama/Llama-3.1-8B-Instruct
Llama-3.2-3B-Instruct 3B meta-llama/Llama-3.2-3B-Instruct
Qwen3-1.7B 1.7B Qwen/Qwen3-1.7B
Qwen3-0.6B 0.6B Qwen/Qwen3-0.6B
@@
Appendix I The details for the fine-tuning PTMs
I.1. Training data collection and processing
Similar to the collection process for the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark data, we gathered an additional 10 annual 10-K financial reports filed with the SEC for the period from February 13, 2024, to February 13, 2025, as summarized in Table 14. These reports contain a total of 33,848 standard taxonomy-tagged facts. Using BeautifulSoup to parse these documents, we identified 22,847 narrative sentences (approximately 5.5 million characters) and 1,236 financial tables. The companies included in this dataset follow the XBRL standard, ensuring comprehensive coverage for training PTMs.
Table 14. Financial report statistics summary for raw training data
@@ I.1. Training data collection and processing
Table 15. Statistics of training data (tokens calculated with “cl100k_base” tokenizer, $\pm$ standard deviation).
@@
However, to align with the extreme classification format used in previous XBRL tagging benchmarks, we directly adopt the US-GAAP tags as entity labels, annotating each token in sentences and tables using the BIO scheme. Specifically, B denotes the beginning of an entity phrase, I marks the continuation (inside) of an entity phrase, and O indicates tokens outside of any entity. As shown in Figure 8, ”4.9” and ”4.5” are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”). To comprehensively cover all US-GAAP tags, we combine the entire set of 17,388 tags from the US-GAAP 2024 taxonomy with the BIO labeling scheme to construct an extreme classification label space, resulting in 34,777 unique entity labels ($2\times 17388+1$).
Figure 8. An example for training set annotation with BIO scheme.
@@
I.2. Training settings
We fine-tune three pretrained models, BERT-large (Devlin et al., 2018), FinBERT (Araci, 2019), and SECBERT (Loukas et al., 2022), on our training set using the HuggingFace Transformers library. All models are trained with a batch size of 4, a learning rate of 3e-5, and for 20 epochs. Optimization is performed using AdamW without gradient accumulation or early stopping. Token classification heads are randomly initialized and trained jointly with the base encoder. Input sequences are tokenized with a maximum length of 512, and labels are aligned at the sub-token level following the BIO tagging scheme. Loss is computed only on the first sub-token of each word to avoid misalignment bias.
@@ I.2. Training settings
Appendix J Retrieval Results in FinCL subtask
@@
Table 16. Acc@k retrieval performance on the FinCL task. FWC: fixed window-based; SAC: structure-aware.
@@
SAC Sentence 0.0696 0.2742 0.3631 0.4798 0.5701 0.6125 0.6535
Table 0.0159 0.0872 0.1245 0.1938 0.2452 0.2723 0.2959
Overall 0.0202 0.1152 0.1534 0.2188 0.2727 0.3012 0.3274
From Table 16, we observe that the SAC strategy consistently outperforms FWC, particularly in the table context, highlighting the importance of aligning the context window with the underlying structural unit.
@@
For sentence-based entities, both strategies yield comparable results, with SAC achieving slightly higher Acc@1 (0.0696 vs. 0.0658) and showing marginal improvements across all cutoff values (e.g., Acc@100 of 0.5701 vs. 0.5618). This suggests that even for relatively unstructured text, preserving sentence boundaries provides minor benefits over fixed-length windows.
@@
In contrast, for table-based entities, the performance gap is substantial. SAC achieves significantly higher retrieval accuracy (e.g., Acc@100 of 0.2452 vs. 0.0050), indicating that row-level context is far more informative than arbitrary character windows when dealing with tabular structures. FWC performs poorly in this setting, likely due to the fragmented and semantically sparse nature of partial table text.
@@
When aggregating results across both structures, SAC outperforms FWC by a wide margin at all retrieval depths (e.g., Acc@200 of 0.3274 vs. 0.0608). These findings underscore the importance of structure-aware context construction, especially in scenarios where inputs span multiple formats such as sentences and tables.
@@
Prompt Template for FinNI Subtask You are a financial information extraction expert specializing in identifying financial numerical entities in XBRL reports. Your task is to extract all such numerical entities from the provided text or serialized <table></table> data and classify them into one of five categories: - "integerItemType": Counts of discrete items, such as the number of employees or total transactions. - "monetaryItemType": Financial amounts expressed in currency, such as revenue, profit, or total assets. - "perShareItemType": Per-share values, such as earnings per share (EPS) or book value per share. - "sharesItemType": Counts of shares, such as outstanding shares or ownership stakes. - "percentItemType": Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol ("%"). Important Instructions: (1) Financial numerical entities are not limited to Arabic numerals (e.g., 10,000). They may also appear in word form (e.g., "ten million"), which must be correctly identified and converted into standard numerical format. (2) Not all numbers in the text should be extracted. Only those that belong to one of the five financial entity categories above should be included. Irrelevant numbers (such as phone numbers, dates, or general IDs) must be ignored. (3) If a number is followed by a magnitude term (e.g., Hundred, Thousand, Million, Billion), do not expand it into the full numerical value. * "Two hundred" -> Extract only "two", not "200". * "10.6 million" -> Extract only "10.6", not "10,600,000". (4) Standardize numerical formatting by removing currency symbols (e.g., "USD"), percentage signs ("%"), and commas (",") while preserving the numeric value. These elements must be removed to ensure consistency. (5) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>}]} Input: {text/table} Output:
Figure 9. Prompt template used for the FinNI subtask.
Prompt Template for FinCL Subtask (Reranking) You are a financial tagging assistant trained in US-GAAP taxonomy. Given a query consisting of an entity, its type, its surrounding context, and the source format (either text or table), your task is to select the single most appropriate US-GAAP tag from a list of 200 candidate tags. Make your decision by carefully analyzing the meaning and context of the entity and matching it with the semantics of the tags. Only output one tag, the best match. Do not explain or list multiple tags. The output is a JSON format: {"result": <the best matched tag>}. Input Query: <entity> + <entity type> + <context> Candidate Tags: {Top 200 US-GAAP tags} Answer:
Figure 10. Prompt template used for the Reranking statge in the FinCL subtask.
@@
Prompt Template for Ablation You are an XBRL tagging expert specializing in annotating financial numerical facts in XBRL reports. Your task is to (1) extract all such numerical entities from the provided text or serialized <table></table> data, (2) classify them into one of five categories, and (3) assign an appropriate US-GAAP tag to each entity. Categories: - "integerItemType": Counts of discrete items, such as the number of employees or total transactions. - "monetaryItemType": Financial amounts expressed in currency, such as revenue, profit, or total assets. - "perShareItemType": Per-share values, such as earnings per share (EPS) or book value per share. - "sharesItemType": Counts of shares, such as outstanding shares or ownership stakes. - "percentItemType": Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol ("%"). US-GAAP tags: - A US-GAAP tag is a standardized semantic label used in XBRL filings to identify specific financial concepts defined by the U.S. Generally Accepted Accounting Principles (GAAP). Each tag represents a distinct accounting item and enables consistent, machine-readable financial reporting. - Examples: "us-gaap:AssetsCurrentAbstract", "us-gaap:AccruedInsuranceNoncurrent". Important Instructions: (1) Financial numerical entities are not limited to Arabic numerals (e.g., 10,000). They may also appear in word form (e.g., "ten million"), which must be correctly identified and converted into standard numerical format. (2) Not all numbers in the text should be extracted. Only those that belong to one of the five financial entity categories above should be included. Irrelevant numbers (such as phone numbers, dates, or general IDs) must be ignored. (3) If a number is followed by a magnitude term (e.g., Hundred, Thousand, Million, Billion), do not expand it into the full numerical value. * "Two hundred" -> Extract only "two", not "200". * "10.6 million" -> Extract only "10.6", not "10,600,000". (4) Standardize numerical formatting by removing currency symbols (e.g., "USD"), percentage signs ("%"), and commas (",") while preserving the numeric value. These elements must be removed to ensure consistency. (5) You should assign the most appropriate US-GAAP tag to each identified entity based on your internal understanding of the 2024 US-GAAP taxonomy. (6) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>, "Tag": <Assigned US-GAAP tag>}]} Input: {text/table} Output:
Figure 11. Prompt template used for ablation study.